The energy required to break a molecular bond in a one mole sample is 965 kJ/mol. What is the longest wavelength of light capable of breaking this bond?

A. 2.06×10-22 nm
B. 1.24×10-7 nm
C. 1.24×10-16 nm
D. 124 nm
E. 1.24×105 nm
1
Expert's answer
2019-04-11T06:00:31-0400

The energy required to break a molecular bond in a one mole sample is 965 kJ/mol. What is the longest wavelength of light capable of breaking this bond?


A. 2.06×10-22 nm

B. 1.24×10-7 nm

C. 1.24×10-16 nm

D. 124 nm

E. 1.24×105 nm


Solution:

ΔH = 965 kJ/mol

c = 3*108 m/s

h = 6.626*10-34

Na = 6.02*1023 mol-1

 

Answer:

λ = 1.24*10-7 nm.